# Refined asymptotics for the Cauchy problem for the fast $p$-Laplace evolution equation

### v2

@@ 6.1 Information on related radial classical and weighted Fast Diffusion Equations
which proves that $\Phi(t,\varrho)=0$ for all $t\geq T$ and $\varrho\geq 0$. Inequality (6.25) plays the role of inequality (6.17) in this range of parameters. Indeed, from (6.25) one can deduce (6.18) for any $0<t<T$ which is enough to establish that $\overline{u}$ is well defined also in the present case. At the same time, using again [16, Lemma 11], we have that $\Phi\in C^{\infty}(0,\infty)^{2}$ and inequalities from (6.19) hold also in the present case. This is enough to show that $\overline{u}$ is a weak solution to (6.20). Moreover, using the same argument as above, one can easily prove inequality (6.21), which is the missing condition to verify in order to use the uniqueness result of [42, Theorem II.1]. We have explained in Section 3.3 that in our setting for a solution $u$ to (CPLE), the function $(t,x)\mapsto\nabla u(t,x)\in C^{\alpha}((0,\infty)\times\mathbb{R}^{N})$. Hence, we are in the position to guarantee that $\partial_{r}u(t,0)=0$ (which means that $u$ is also a solution to (6.20)) and therefore, by the uniqueness, we have that $u=\overline{u}$. This concludes the proof in the case $p<p_{c}$.
@@
To the best of our knowledge, the problem (6.20) has not been investigated yet. It seems, as well, that the Neumann problem for $p$-Laplacian type equation has been much less studied. For more information we refer to [6, 5] in the case of a Neumann problem in bounded domains, to [4] for the Neumann problem for the Porous Medium Equation ($\partial_{t}u=\Delta u^{m}$, $m>1$), and to [71, Chapter 11] for exposition of the background in detail. In dimension $N=1$, the techniques used in the seminal paper [43] can be adapted (at least in the good range $p_{c}<p<2$) in order to prove the existence, uniqueness and comparison principle. We also stress that problem (6.20) is very similar to (CPLE) and the techniques of [42] can be adapted in the whole generality for the entire range $21<p<$ 2.
@@ 6.2 Proof of the convergence in the relative error of the radial derivatives
$\begin{cases}\partial_{\tau}\Psi=|y|^{{\mathfrak{a}}}\,\mathrm{div}\big{[}|y|^{-{\mathfrak{a}}}\,\left(\nabla\Psi^{m}-y\,\Psi\right)\big{]}\,\quad\text{for }\ (\tau,y)\in\left(0,\infty\right)\times\mathbb{R}^{N}\,,\\ \Psi(0,y)=\Psi_{0}(|y|)\,\quad\text{for }\ y\in\mathbb{R}^{N}\,,\end{cases}$ (6.27)
@@
$-\varepsilon(t)\,\mathfrak{B}_{\overline{D},T}(t,\varrho)\leq\Phi(t,\varrho)-\mathfrak{B}_{\overline{D},T}(t,\varrho)\leq\varepsilon(t)\,\mathfrak{B}_{\overline{D},T}(t,\varrho)\quad\mbox{where}\quad\varepsilon(t)=C_{\diamond}\,(T-t)^{-\lambda}\,.$
@@
$-\varepsilon(t)\,\mathcal{B}_{D,T}(t,r)\leq u(t,r)-\mathcal{B}_{D,T}(t,r)\leq\varepsilon(t)\,\mathcal{B}_{D,T}(t,r)\quad\forall\,r\geq 0\quad\forall\,\,t_{\diamond}<t<T\,,$
@@ 7 Proof of Theorem 2 and related results
Let $N\geq 1$, $21<p<$ 2, $0\leq v\in C^{1,\alpha}(\mathbb{R}^{N})$ for some $10<\alpha<$ 1, and $D>0$ such that $v$ satisfies (A1) for some ${\varepsilon}\in\left(0,1\right)$ and
@@ 8 Proofs of Theorem 4 and 5 capturing in particular the subcritical case $p\leq p_{c}$
Step 1): identification of the limit when $\tau\rightarrow\infty$. For this step, we follow mainly [2, Lemma 2.5]. We will first prove that $v(\tau,\cdot)$ converges to $V_{D}$ pointwise and in $\mathrm{L}^{p}$-norms. As in [2, Lemma 2.5], let us define $v^{h}(\tau,y):=v(\tau+h,y)$, for any given $h>0$ and $\tau\in\left[0,1\right]$. By the comparison principle and thanks to assumption (i), $\{v^{h}\}$ is uniformly bounded. Furthermore, it is uniformly continuous in $[0,1]\times B_{R}$ thanks to assumption (iv). By the Ascoli–Arzelá Theorem, for any sequence $h_{n}\rightarrow\infty$ (as $n\rightarrow\infty$) the sequence of functions $\{v^{h_{n}}\}$ converges uniformly (up to a subsequence) to a function $v^{\infty}$ on compact subsets of $[0,1]\times{\mathbb{R}^{N}}$. Moreover, we infer that for any $R>0$ it holds that $\|v\|_{C^{1,\alpha}([0,1]\times B_{R})}<\infty$ and for $\tau\in\left[0,1\right]$ function $v^{\infty}(\tau,\cdot)$ satisfies (i). Since $N<\frac{p}{(2-p)(p-1)}$, we know that $V_{D}^{\gamma-2}(v-V_{D})^{2}\in\mathrm{L}^{1}({\mathbb{R}^{N}})$, cf. (5.1) and the end of the proof of Lemma 5.7. Therefore, by using the arguments of Lemma 5.6, we get that ${\mathsf{E}}[v],{\mathcal{E}}[v|{V_{D}}]<\infty$. Thanks to assumption (iii) the entropy functional ${\mathcal{E}}[v|{V_{D}}]$ is non-negative and $\tau\mapsto{\mathcal{E}}[v(\tau)|{V_{D}}]$ is decreasing in time. By the time monotonicity, ${\mathcal{E}}[v^{h_{n}}(\tau)|{V_{D}}]$ and ${\mathcal{E}}[v^{h_{n}+1}(\tau)|{V_{D}}]$ have the same limit for $h_{n}\rightarrow\infty$. Therefore we infer that
@@
By the positivity of ${\mathcal{I}}[v^{\infty}(\tau)|{V_{D}}]$ and Fatou’s Lemma, we infer therefore that $00\geq\int_{0}^{1}{\mathcal{I}}[v^{\infty}(\tau)|{V_{D}}]\,{\rm d}\tau=$ 0. Consequently, $\nabla(v^{\infty})^{\gamma-1}=\nabla V_{D^{\star}}^{\gamma-1}$ for some $D^{\star}>0$ and so $v^{\infty}=V_{D^{\star}}$. Up to now, we have proven that $v^{h_{n}}$ converges pointwise towards $V_{D^{\star}}$ as $n\rightarrow\infty$. We only need to ensure that $D^{\star}=D$. By Lemma A.2, we are in the position of using the Dominated Convergence Theorem to infer that $\{(v^{h_{n}}-V_{D^{\star}})\}$ converges in $\mathrm{L}^{1}({\mathbb{R}^{N}})$, which implies that
@@
[…]
@@
where $C(\alpha,\tau_{0},{{v_{0}}})$ is as in (iv). Combining the above estimate with (8.5) one get (8.6) with $\tau_{\star}=\max\{\tau_{0},-\frac{2}{\mathfrak{K}}\log(H\,\varepsilon^{\frac{3-2p}{2-p}})\}$, for a constant $H=H(\alpha,\tau_{0},{{v_{0}}})>0$. We notice that for ${\varepsilon}$ small enough we have that $-\frac{2}{\mathfrak{K}}\log(H\,\varepsilon^{\frac{3-2p}{2-p}})>\max\{0,\tau_{0}\}$. Once inequality (8.6) is obtained with an explicit functional relation between ${\varepsilon}$ and $\tau_{\star}({\varepsilon})$, one can compute the rate of convergence by inverting this relation, as it was done, for instance, in [18, Corollary 4.14]. This is enough to obtain the convergence result (1.34) for the $\mathrm{L}^{\infty}$-norm. The result in the $\mathrm{L}^{q}$-norm is obtained by interpolation as it is done in the proof of Theorem 4. The proof is complete. ∎
@@ 9 Justification of the assumption for Theorems 4 and 5 in the radial decreasing case
where $r=|y|$. Let us consider $\psi\mathbb{R}^{N}\to\left[0,\infty\right)$ a radial cut-off function $\psi(y)=\psi(r)$ which is equal to $1$ when $0\leq r\leq 1$ and equal to 00 when $r\geq 2$. Let us take $\phi(y)=\psi(y/R)$ in (9.3). Since $|\nabla\psi|\leq c\,R^{-1}$, where $c>0$ is a dimensionless constant. Since $\partial_{r}\psi(r)\neq 0$ only when $R<r<2R$ we find, by applying Lemma A.2, that
@@
since $11+(\gamma-2)(p-2)+\frac{p-2}{p-1}=$ 1. Under the current assumptions, in a neighbourhood of the origin, we can bound the term $|\nabla v^{\gamma-1}(\tau,y)-\nabla V_{D}^{\gamma-1}(y)|^{2}$ as
@@
By idenity (6.3), i.e. $-\partial_{r}u(t,r)={\mathcal{D}}\,{\varrho}^{\frac{2}{m+1}}\,\Phi(t,{\varrho})$, we have that $\Phi:\mathbb{R}^{N}\times(0,\infty)\rightarrow\mathbb{R}$) is a nonnegative radial function of the variables $({\varrho}=|x|,t)$ which solves, by Theorem 12, the following weighted equation
$\partial_{t}\Phi={\varrho}^{1-n}\,\partial_{{\varrho}}\left({\varrho}^{n-1}\,\partial_{{\varrho}}\Phi^{m}\right)\,,\qquad m=p-1\,.$
We notice that Claim 2 follows essentially from the comparison principle (i.e. property (3.3)) for equation (R-CPLE).
In order to prove Claim 2, there are two cases to be considered: the simpler one is when $n=2+2N\frac{(p-1)}{p}$ (which is always bigger than 2) is an integer, and in this case $\Phi$ is the radial solution to a FDE with $m=p-1$ with the classical Laplacian in dimension $n$, and the claim follows by [10, Theorem 4]. However, this does not happen in general, hence we introduce an operator with suitable CKN-type weights: consider $\Phi$ as a radial function of $R^{N}$, then we have, according to (6.5), that
@@
$\int(\Phi_{1}(t,x)-\Phi_{2}(t,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x\leq\int(\Phi_{1}(0,x)-\Phi_{2}(0,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x\qquad\mbox{for all }t>0\,.$
for a suitable positive and radially decreasing $\psi$, such that $\mathcal{L}_{\mathfrak{a}}\psi\leq 0$. We shall see an example of it below. The above inequality easily implies that if $\Phi_{1}(0,\cdot)\leq\Phi_{2}(0,\cdot)$ then $\Phi_{1}(t,\cdot)\leq\Phi_{2}(t,\cdot)$ for all $t>0$. Undoing the change of functions and going back to PLE variables, this is exactly Claim 2.
@@
$\psi(r)=(1+r^{2})^{-\frac{n-2}{2}}\qquad\mbox{is such that}\qquad\mathcal{L}_{\mathfrak{a}}\psi=-n(n-2)\psi^{\frac{2n}{n-2}}\leq 0$
We present now a proof that holds for strong solutions [^1] For mild solutions, i.e. nonlinear gradient flows on the Banach space $L^{1}(|x|^{-{\mathfrak{a}}}\,{\rm d}x)$, this property is often called $T$-Contraction or “well ordering” of the nonlinear semigroup. This always holds for mild solutions, see for instance [72] and references therein. Strong solutions are particular cases of mild solutions, hence the result holds. We have decided to sketch the proof, since it gives an idea of how it works for weak solution: this is done by a careful choice of test function in the weak formulation: take a smooth, compactly supported $\varphi_{i,j,k}(\tilde{t},x)=\eta_{i}(\tilde{t})\psi_{k}(|x|)\zeta_{j}(\tilde{t},x)$, where $\eta_{i}(\tilde{t})\xrightarrow[]{i\to\infty}\chi_{[}0,t](\tilde{t})$ and $\psi_{k}\xrightarrow[]{k\to\infty}\psi$ and with some additional properties, cf. Appendix of [22]. We also take a smooth approximation $\zeta_{i}(\tilde{t},x)\xrightarrow[]{j\to\infty}H(\Phi_{1}(\tilde{t},x)-\Phi_{2}(\tilde{t},x))$., i.e. for those solutions such that the equation holds almost everywhere, i.e. for which $\partial_{t}\Phi=\mathcal{L}_{\mathfrak{a}}\Phi^{m}\in\mathrm{L}^{1}_{\rm loc}((0,\infty)\times\mathbb{R}^{N})$. Notice that for strong solutions we always have that
$\begin{split}\int(\Phi_{1}(t,x)-\Phi_{2}(t,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x&-\int(\Phi_{1}(0,x)-\Phi_{2}(0,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x\\ &=\int_{0}^{t}\frac{{\rm d}}{{\rm d}\tilde{t}}\int(\Phi_{1}(\tilde{t},x)-\Phi_{2}(\tilde{t},x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x{\rm d}\tilde{t}\,,\end{split}$
@@
It only remains to be proven the validity of condition $(iv)$ and inequality (1.33) of Theorem 5 under the additional hypothesis (9.1).
Proof that condition $(iv)$ of Theorem 4 holds under the additional hypothesis (9.1). The proof of the uniform regularity estimates for radial functions follows by extending regularity estimates for the FDE with CKN-weights to the present (easier since radial) case: here for some values of ${\mathfrak{a}}$ we fall out the “classical” CKN-setting of [22, 23], since a priori we do not have the corresponding CKN inequalities that allow to perform the Nash iteration. However, since we have opposite powers inside and outside the divergence, and we are in the radial setting, we luckily have “the right” weighted Sobolev type inequality valid for all $n>2$ that reads:

### v3

@@ 6.1 Information on related radial classical and weighted Fast Diffusion Equations
which proves that $\Phi(t,\varrho)=0$ for all $t\geq T$ and $\varrho\geq 0$. Inequality (6.25) plays the role of inequality (6.17) in this range of parameters. Indeed, from (6.25) one can deduce (6.18) for any $0<t<T$ which is enough to establish that $\overline{u}$ is well defined also in the present case. At the same time, using again [16, Lemma 11], we have that $\Phi\in C^{\infty}(0,\infty)^{2}$ and inequalities from (6.19) hold also in the present case. This is enough to show that $\overline{u}$ is a weak solution to (6.20). Moreover, using the same argument as above, one can easily prove inequality (6.21), which is the missing condition to verify in order to use the uniqueness result of [43, Theorem II.1]. We have explained in Section 3.3 that in our setting for a solution $u$ to (CPLE), the function $(t,x)\mapsto\nabla u(t,x)\in C^{\alpha}((0,\infty)\times\mathbb{R}^{N})$. Hence, we are in the position to guarantee that $\partial_{r}u(t,0)=0$ (which means that $u$ is also a solution to (6.20)) and therefore, by the uniqueness, we have that $u=\overline{u}$. This concludes the proof in the case $p<p_{c}$.
@@
To the best of our knowledge, the problem (6.20) has not been investigated yet. It seems, as well, that the Neumann problem for $p$-Laplacian type equation has been much less studied. For more information we refer to [6, 5] in the case of a Neumann problem in bounded domains, to [4] for the Neumann problem for the Porous Medium Equation ($\partial_{t}u=\Delta u^{m}$, $m>1$), and to [72, Chapter 11] for exposition of the background in detail. In dimension $N=1$, the techniques used in the seminal paper [44] can be adapted (at least in the good range $p_{c}<p<2$) in order to prove the existence, uniqueness and comparison principle. We also stress that problem (6.20) is very similar to (CPLE) and the techniques of [43] can be adapted in the whole generality for the entire range $1<p<2$.
@@ 6.2 Proof of the convergence in the relative error of the radial derivatives
[…]
@@
$\varepsilon(t)\,\mathfrak{B}_{\overline{D},T}(t,\varrho)\leq\Phi(t,\varrho)-\mathfrak{B}_{\overline{D},T}(t,\varrho)\leq\varepsilon(t)\,\mathfrak{B}_{\overline{D},T}(t,\varrho)\quad\mbox{where}\quad\varepsilon(t)=C_{\diamond}\,(T-t)^{-\lambda}\,.$
@@
$\varepsilon(t)\,\mathcal{B}_{D,T}(t,r)\leq u(t,r)-\mathcal{B}_{D,T}(t,r)\leq\varepsilon(t)\,\mathcal{B}_{D,T}(t,r)\quad\forall\,r\geq 0\quad\forall\,\,t_{\diamond}<t<T\,$,
@@ 7 Proof of Theorem 2 and related results
Let $N\geq 1$, $1<p<2$, $0\leq v\in C^{1,\alpha}(\mathbb{R}^{N})$ for some $0<\alpha<1$, and $D>0$ such that $v$ satisfies (A1) for some ${\varepsilon}\in\left(0,1\right)$ and
@@ 8 Proofs of Theorem 4 and 5 capturing in particular the subcritical case $p\leq p_{c}$
Step 1): identification of the limit when $\tau\rightarrow\infty$. For this step, we follow mainly [2, Lemma 2.5]. We will first prove that $v(\tau,\cdot)$ converges to $V_{D}$ pointwise and in $\mathrm{L}^{p}$-norms. As in [2, Lemma 2.5], let us define $v^{h}(\tau,y):=v(\tau+h,y)$, for any given $h>0$ and $\tau\in\left[0,1\right]$. By the comparison principle and thanks to assumption (i), $\{v^{h}\}$ is uniformly bounded. Furthermore, it is uniformly continuous in $0,1]\times B_{R}$ thanks to assumption (iv). By the Ascoli–Arzelá Theorem, for any sequence $h_{n}\rightarrow\infty$ (as $n\rightarrow\infty$) the sequence of functions $\{v^{h_{n}}\}$ converges uniformly (up to a subsequence) to a function $v^{\infty}$ on compact subsets of $0,1]\times{\mathbb{R}^{N}}$. Moreover, we infer that for any $R>0$ it holds that $\|v\|_{C^{1,\alpha}([0,1]\times B_{R})}<\infty$ and for $\tau\in\left[0,1\right]$ function $v^{\infty}(\tau,\cdot)$ satisfies (i). Since $N<\frac{p}{(2-p)(p-1)}$, we know that $V_{D}^{\gamma-2}(v-V_{D})^{2}\in\mathrm{L}^{1}({\mathbb{R}^{N}})$, cf. (5.1) and the end of the proof of Lemma 5.7. Therefore, by using the arguments of Lemma 5.6, we get that ${\mathsf{E}}[v],{\mathcal{E}}[v|{V_{D}}]<\infty$. Thanks to assumption (iii) the entropy functional ${\mathcal{E}}[v|{V_{D}}]$ is non-negative and $\tau\mapsto{\mathcal{E}}[v(\tau)|{V_{D}}]$ is decreasing in time. By the time monotonicity, ${\mathcal{E}}[v^{h_{n}}(\tau)|{V_{D}}]$ and ${\mathcal{E}}[v^{h_{n}+1}(\tau)|{V_{D}}]$ have the same limit for $h_{n}\rightarrow\infty$. Therefore we infer that
@@
By the positivity of ${\mathcal{I}}[v^{\infty}(\tau)|{V_{D}}]$ and Fatou’s Lemma, we infer therefore that $0\geq\int_{0}^{1}{\mathcal{I}}[v^{\infty}(\tau)|{V_{D}}]\,{\rm d}\tau=0$. Consequently, $\nabla(v^{\infty})^{\gamma-1}=\nabla V_{D^{\star}}^{\gamma-1}$ for some $D^{\star}>0$ and so $v^{\infty}=V_{D^{\star}}$. Up to now, we have proven that $v^{h_{n}}$ converges pointwise towards $V_{D^{\star}}$ as $n\rightarrow\infty$. We only need to ensure that $D^{\star}=D$. By Lemma A.2, we are in the position of using the Dominated Convergence Theorem to infer that $\{(v^{h_{n}}-V_{D^{\star}})\}$ converges in $\mathrm{L}^{1}({\mathbb{R}^{N}})$, which implies that
@@
$\text{$\kappa_{1}\,R_{\varepsilon}^{\frac{-p}{p-1}}<\varepsilon\quad$ and $\quad\kappa_{2}\,\|v(\tau,y)-V_{D}(y)\|_{\mathrm{L}^{\infty}({\mathbb{R}^{N}})}\,R_{\varepsilon}^{\frac{p}{2-p}}<\varepsilon$}\,.$
@@
where $C(\alpha,\tau_{0},{{v_{0}}})$ is as in (iv). Combining the above estimate with (8.5) one get (8.6) with $\tau_{\star}=\max\{\tau_{0},-\frac{2}{\mathfrak{K}}\log(H\,\varepsilon^{\frac{3-2p}{2-p}})\}$, for a constant $H=H(\alpha,\tau_{0},{{v_{0}}})>0$. We notice that for $\varepsilon$} small enough we have that $\frac{2}{\mathfrak{K}}\log(H\,\varepsilon^{\frac{3-2p}{2-p}})>\max\{0,\tau_{0}\}$. Once inequality (8.6) is obtained with an explicit functional relation between $\varepsilon$} and $\tau_{\star}({\varepsilon})$, one can compute the rate of convergence by inverting this relation, as it was done, for instance, in [18, Corollary 4.14]. This is enough to obtain the convergence result (1.34) for the $\mathrm{L}^{\infty}$-norm. The result in the $\mathrm{L}^{q}$-norm is obtained by interpolation as it is done in the proof of Theorem 4. The proof is complete. ∎
@@ 9 Justification of the assumption for Theorems 4 and 5 in the radial decreasing case
where $r=|y|$. Let us consider $\psi\mathbb{R}^{N}\to\left[0,\infty\right)$ a radial cut-off function $\psi(y)=\psi(r)$ which is equal to $1$ when $0\leq r\leq 1$ and equal to 0 when $r\geq 2$. Let us take $\phi(y)=\psi(y/R)$ in (9.3). Since $|\nabla\psi|\leq c\,R^{-1}$, where $c>0$ is a dimensionless constant. Since $\partial_{r}\psi(r)\neq 0$ only when $R<r<2R$ we find, by applying Lemma A.2, that
@@
since $1+(\gamma-2)(p-2)+\frac{p-2}{p-1}=1$. Under the current assumptions, in a neighbourhood of the origin, we can bound the term $|\nabla v^{\gamma-1}(\tau,y)-\nabla V_{D}^{\gamma-1}(y)|^{2}$ as
@@
By idenity (6.3), i.e. $\partial_{r}u(t,r)={\mathcal{D}}\,{\varrho}^{\frac{2}{m+1}}\,\Phi(t,{\varrho})$, we have that $\Phi:\mathbb{R}^{N}\times(0,\infty)\rightarrow\mathbb{R}$) is a nonnegative radial function of the variables $({\varrho}=|x|,t)$ which solves, by Theorem 12, the following weighted equation
$\partial_{t}\Phi={\varrho}^{1-n}\,\partial_{{\varrho}}\left({\varrho}^{n-1}\,\partial_{{\varrho}}\Phi^{m}\right)\,,\qquad m=p-1\,.$ (9.8)
We notice that Claim 2 follows essentially from the comparison principle (i.e. property (3.3)) for derivatives of equation (R-CPLE). As explained in Section 6, radial derivatives of a solution to (R-CPLE) (up to a power of $\rho$) satisfy equation (6.27), which is a rescaled version of (9.8). We concluded therefore that, in order to prove Claim 2, it suffices to prove a comparison principle for equation (9.8). There are two cases to be considered: the simpler one is when $n=2+2N\frac{(p-1)}{p}$ (which is always bigger than 2) is an integer, and in this case $\Phi$ can be seen as the radial solution to a FDE with m=p1m=p-1 with the classical Laplacian in dimension $n$, and the claim follows by [10, Theorem 4]. However, this does not happen in general, hence we introduce an operator with suitable CKN-type weights: consider $\Phi$ as a radial function of $R^{N}$, then we have, according to (6.5), that
@@
$\int(\Phi_{1}(t,x)-\Phi_{2}(t,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x\leq\int(\Phi_{1}(0,x)-\Phi_{2}(0,x))_{+}\psi(|x|)|x|^{-{\mathfrak{a}}}\,{\rm d}x\qquad\mbox{for all }t>0\,.$ (9.11)
for a suitable positive and radially decreasing $\psi$, such that $\mathcal{L}_{\mathfrak{a}}\psi\leq 0$. We shall see an example of it below. The above inequality easily implies that if $\Phi_{1}(0,\cdot)\leq\Phi_{2}(0,\cdot)$ then $\Phi_{1}(t,\cdot)\leq\Phi_{2}(t,\cdot)$ for all $t>0$. Claim 2 follows from this last step, as explained above.
@@
$\psi(r)=(1+r^{2})^{-\frac{n-2}{2}}\qquad\mbox{is such that}\qquad\mathcal{L}_{\mathfrak{a}}\psi=-n(n-2)\psi^{\frac{n+2}{n-2}}\leq 0$
[…]
@@
It only remains to prove the validity of condition $(iv)$ and inequality (1.33) of Theorem 5 under the additional hypothesis (9.1).
Proof that condition $(iv)$ of Theorem 4 holds under the additional hypothesis (9.1). The proof of the uniform regularity estimates for radial functions follows by extending regularity estimates for the FDE with CKN-weights to the present (easier since radial) case: here for some values of $\mathfrak{a}$} we fall out the “classical” CKN-setting of [22, 23], since a priori we do not have the corresponding CKN inequalities that allow to perform the Nash-Moser iteration. However, since we have opposite powers inside and outside the divergence, and we are in the radial setting, we luckily have “the right” weighted Sobolev type inequality valid for all $n>2$ that reads: